An electric bulb, when connected across a power supply of 220 V, consumes a power of 60 W. If the supply drops to 180 V, what will be the power consumed? If the supply is suddenly increased to 240 V, what will be the power consumed?

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. E = 2220 v P = 60 w R = V^2/P = 220 * 220/60 = 220 * 11/3 Ω (a) E = 180 v P = V^2/R = 180 * 180 * 3/220 *11 = 40.16 = 40 w (b) E = 240 v P = V^2/R = 240 *240 *3/220 *11 = 71.4 = 71 w